A landscaper, who just completed a rectangular flower garden measuring 6 feet by 10 feet, orders 1 cubic yard of premixed cement, all of which is to be used to create a border of uniform width around the garden. If the border is to have a depth of 3 inches, how wide will the border be?
Found 2 solutions by mananth, dharvin1965@yahoo.com:
Answer by mananth(16949)

garden = 6ft by 10ft. Area of garden = 60 sq. ft.
1 cu. yard = 3*3*3 = 27 cu.ft
..
The volume of concrete border = 27cu.ft.
the thickness = 3 inches. = 3/12 ft = 1/4 ft.
...
volume of concrete = L*W*H
L*W= volume / height
volume = 27cu.ft
height or thickness = 3in.
Area of the border = 27/(1/4 )
L*W= 27*4/1 = 108 sq. inches.
..
area of garden + area of border = 60+108 = 168 sq.ft
let width of border be x feet. increase in length & width =2x
(6+2x)(10+2x)= 168 sq. ft
60+12x+20x+4x^2=168
4x^2+32x+60-168 =0
4x^2+32x-108=0
/4
x^2+8x-27 =0
solve by quadratic formula you get x = 2.56 feet.
The width will be 2.56 feet

Answer by dharvin1965@yahoo.com(1)

arectangularflowergardenhasanereaof168squarefeetifthewidthofthegardenis12feetthenwhatwhatisthelengthofthegarden